OpenStudy (anonymous):

Find the area of the triangle with A = 60°, b = 11 feet, and c = 8 feet. Round to the nearest tenth. A. 22 ft.^2 B. 38.1 ft.^2 C. 14.1 ft.^2 D. 12.4 ft.^2

OpenStudy (jmark):

Given the area of the triangle with A = 60°, b = 11 feet, and c = 8 feet. a^2=b^2+c^2- 2bccosA // to find the value of a // a^2 = 121 + 64 - 176 cos 60 a = √97 a = 9.8 feet s = (a + b + c)/2 = (9.8 + 11 + 8)/2 = 14.4 A = √s(s - a)(s - b)(s – c) // area of triangle // = √[14.4(14.4 - 9.8)(14.4 - 11)(14.4 - 8)] A = 37.96 = 38 ft^2 Learn more about Area of Circle http://goo.gl/dbXhCw